OpenStudy (anonymous):

Find the volume of a cone with radius 28 and height 44. Answer 36124.13 31246.13 34612.13 36421.13

OpenStudy (anonymous):

I got 36124.13 the first one haha.

OpenStudy (anonymous):

but im not sure if its right?

OpenStudy (austinl):

Do the same as the last problem. \[V = \frac{1}{3}\times \pi\times r^2\times h\]

OpenStudy (anonymous):

it gave me 36124.13

OpenStudy (austinl):

\[V = \frac{1}{3}\times \pi\times 28^2\times 44=36,124.1\]

OpenStudy (austinl):

That should be your answer.

OpenStudy (anonymous):

okay thx.! :)

OpenStudy (austinl):

No problemo!